# Autonomous Vehicles Using Multi-Agent Reinforcement Learning for Routing Decisions Can Harm Urban Traffic

### v1

@@ 3 Methodology
Figure 2: We model the routing ”game” between humans and AVs on urban traffic networks using RouteRL (Psarou et al., 2024), a custom PettingZoo (Terry et al., 2021) environment, communicating with SUMO as it trains optimal routing policies with standard MARL algorithms via torchRL (Bou et al., 2023). Every day, agents (humans or/and AVs) select a route and their actions are fed into SUMO to generate travel times (rewards), which are then used to update the policy for the next episode (day).
4 Experiments
Figure 3: Single AV replicates human choices, and each tested RL algorithm finds the optimal solution to the binary choice problem. During the training, humans are barely affected by suboptimal actions, and when trained, the system is back in its original equilibrium state. The line represents the mean across realizations and the error band accounts for variations in agent start times and application of different algorithms, as described in Section 4.2.
@@ 4 Experiments
We consider two plausible equilibria resulting from human collective actions: first, when all humans select the shorter route ([$[\{0\}^{22}]$, Figure 1b) and second when they all opt for the longer route ([$[\{1\}^{22}]$, Figure 1c). Both meet Nash criteria for User Equilibrium (Wardrop, 1952) (common paraphrase of Nash equilibrium for the route choice context), with the latter being also System Optimal (minimizing total travel time in the system (Merchant & Nemhauser, 1978)). None of the drivers is inclined to change their route, as it would reduce individual rewards (travel longer and arrive later). For stability, we fix the route for the first two agents to stabilize the early loading of the system. After 200 days of simulation, the system is stable (the next day the rational drivers will replicate today’s choices), fair (travel times are equal among drivers), and either globally optimal (Figure 1b where total travel costs are minimized) or suboptimal (Figure 1c). The individual and system costs (travel times) are reported in Table 1.
@@
In the equilibrated human system, we first replace a random human vehicle with an RL-controlled AV. Its traffic properties (reaction, acceptance gap, acceleration profile, and other IDM parameters (Treiber et al., 2000)) remain intact. AVs are indistinguishable from humans by all but routing decisions: they may use any RL (MARL) algorithm to converge to the optimal policy.
We demonstrate that RL quickly finds an optimal solution, and AV agents follow the crowd (replicate the decision of the human that drove that vehicle before). Unsurprisingly, since the problem is a trivial binary decision in a static environment, this is true for all suitable RL algorithms, including Deep Q-Learning (DQN), PPO, and SAC as we demonstrate in Figure 3. In any equilibrated system, any RL algorithm training AVs with the same reward formulation, and action space as their human predecessors will replicate the optimal strategy, which can be derived directly from conditions of Nash equilibrium. Moreover, in a dense traffic environment, the impact of a single vehicle is unlikely to be noticed by other humans. The marginal cost of actions taken by a single AV to other humans will fall within what is known as the indifference band (Di et al., 2017), making them indistinguishable from the traffic stochastic noise (Neun et al., 2023). Our position, that AVs will disequilibrate the traffic networks, does not hold for a single AV in a static environment, which will converge to a solution that neither deviates from its human predecessor nor impacts other human drivers.
Figure 4: 8000 episodes (that correspond to over 20 days of commute) of training for selected MARL algorithms. The trivial solution is found only by IPPO and MAPPO, other algorithms fail to converge to optimal policy. Suboptimal equilibrium is reproduced after 1000 days but System optimal requires much more training, far exceeding the patience threshold for the remaining 12 human drivers in the system. The first 200 episodes represent the human learning phase, followed by the training and testing phases of AVs. See Appendix B for the experiments’ setup.
@@
Already in our simple scenario, serious issues arise when multiple vehicles learn optimal policies simultaneously. Most likely, more than one AV will be introduced into our cities (whenever they become ready to operate reliably), each solving the same problem of identifying optimal routing policies to reach its destination. Such a multi-agent setting allows a significant alteration to the initial problem: the AVs may communicate (becoming the so-called CAVs) and share data.
Figure 5: With an increasing number of agents, a fraction of those, who learn optimal policy decreases, achieving similar efficiency to random choice (route 0: system optimal - user equilibrium, route 1: suboptimal user equilibrium). This appears to happen with as few as four agents for some of the tested algorithms (ISAC, IQL) that failed to converge in Figure 4. See Appendix B for the experiments’ setup.
First, we demonstrate the natural first stage of AV introduction: with no communication and reward formulation identical to the one of the selfish human drivers. Now, the environment has become non-stationary (Jiang et al., 2024), as the state transitions and rewards of an agent are influenced by the evolving policies of other agents. With as few as 10 AVs (where non-convergence appears after the introduction of just 4 AV agents discussed in Section 4.5 and Appendix G) our setting is sufficient to argue for our position.
@@
In Figure 4, we present two arguments supporting our position. First is the class of algorithms that failed to converge after sufficiently long training (QMIX, VDN, IQL, MASAC, ISAC). Specifically, the choices are far from optimal, as reflected in the noisy profiles that fluctuate around 0.5. QMIX and VDN probably do not find optimal solutions due to the cooperative nature of their implementation. The difficulty of achieving stability and convergence in ISAC and MASAC can be attributed to their off-policy nature combined with the nonlinear function approximation used in neural networks (Haarnoja et al., 2018b; Maei et al., 2009). In such a setting, the consequences of introducing MARL routing algorithms to AVs are negative to all parties: not only the travel times are longer for all agents (humans and AVs) but they also become variable, as we report in Table 1. IPPO and MAPPO, however, converged after a lengthy training.
Figure 6: Top. IPPO and MAPPO still converge if the adaptation term is added to the model (dotted lines). However, the system diverts from the optimal state to the suboptimal, where the total costs are higher. Bottom. The centralized version of IPPO can, in some cases, accelerate convergence, reducing othe time required to reach an optimal solution. However, this reduction is limited to a few days. See Appendix B for the experiments’ setup.
@@
Some MARL algorithms converged and successfully managed to identify optimal policies (the small joint action space of 1024 can be easily enumerated to explicitly identify optimal solutions). Eventually, the 10 AVs managed to stabilize their actions and reliably make optimal choices (Figure 4). After that, the negative impact on the system and the humans diminishes (Table 1). Nonetheless, the learning process was lengthy, with thousands of episodes (days).
@@
The training can be interpreted twofold: within a virtual environment or in a real system. The former is neutral to the real system and its users since AVs can train their policies virtually and deploy them only after training is complete and the policies have converged to optimal solutions. Humans will not be affected and iterations remain only virtual. This, however, requires virtual environment suitable to train a policy, which, is not available to date, as we argued in Section 2.4. If so, the learning period needs to be treated physically and algorithmic iterations are not abstract episodes anymore, but physical days of real disturbances. Eventually, disturbances diminish (as for IPPO and MAPPO), yet the negative impact on the system and its users accumulates (to values presented in Appendix D). Alternatively, RL can be inaccurately trained on imperfect models and collected data and later fine-tuned in reality, as demonstrated in Nair et al. (2021). However, sim2real transfer presents significant challenges, as discussed in Zhao et al. (2020). Real-world complexity exceeds any simulation’s capabilities.
@@
As demonstrated, the problem behind our position is in multiple simultaneously learning agents. To find how many agents can learn at the same time without negatively affecting the system, we did simulations with a gradually increasing number of agents and reported when convergence issues arose. In each simulation, the specific AVs were chosen at random, with the condition that no two AVs are consecutive (there is always a human agent between AVs), and the first two vehicles never mutate. The algorithms start failing to converge even with 3-5 agents, as shown in Figure 5 and Appendix G. This reveals a systemic issue that becomes increasingly prevalent as the number of agents grows, which will pose a serious threat to traffic performance when more AVs participate in our daily commute.
4.6 Human adaptation.
As the system gets disequilibrated, humans will naturally seek ways to improve their payoffs (to arrive faster) (Watling & Cantarella, 2013). This behavior is similar to the process of finding the equilibrium (Bie & Lo, 2010), but even less predictable (He & Liu, 2012). This brings another source of non-stationarity to the system. Now, two paradigms are applied to a single system: humans, rational decision-makers naturally adapt and AVs’ algorithms explore and exploit the environment to identify optimal policies. We incorporate a probabilistic adaptation formula (to the human decisions) and observe that the previously optimal system now shifts to the suboptimal state Figure 6 top). Decisions of adapting humans may shift towards suboptimal equilibria during the AV training period - leading to worse performance for all users.
@@
The final aspect of our position lies in the communication, collaboration, and/or centralization (Schwarting et al., 2019) - which makes AVs the CAVs. This allows us to better exploit the potential of autonomous driving, presumably at the cost of sharing private information with others or with central agents (Nayak et al., 2021). Can we trade our private destination and origin to resolve the non-stationarity issues? To some extent, centralization may speed up the convergence, yet nowhere close to solving the problem and leading to issues with combinatorically growing search space (see Figure 6 bottom).
@@ 5 Conclusion
AVs have begun appearing in traffic networks worldwide, and their presence is expected to increase in the upcoming years. As AVs become more prevalent, MARL can be applied to optimize their route choices, and this optimization can improve urban efficiency and reduce congestion and pollution in our cities. In practice, our MARL algorithms need hundreds of episodes to land in optimal policies, even in trivial cases. To realistically simulate human route-choice behavior we miss theories, data to verify them, and simulation tools. Moreover, realistic urban mobility simulators are lacking, and the training episodes would probably need to be deployed directly on real traffic systems, disrupting the traffic networks. To tackle this problem, we are developing a benchmark tailored to test different road networks, and demand patterns to address the MARL challenges identified in this paper. We hope this will spark discussions within the MARL community encouraging the further development of benchmarks and the exploration of MARL algorithms that can effectively enhance urban mobility. Such an experimental research program is needed to ensure we fully exploit the opportunities of the new technology (AVs) and algorithms (MARL) to help us improve the traffic in future cities.
@@ 6 Alternative Views
What if the limitations of the virtual environment are specific to this study and could be resolved with better design or tools?
As discussed in Section 2.4, accurately representing dynamic, multi-agent systems like traffic networks requires highly complex simulation environments. Modeling realistic traffic demand patterns involves acquiring and handling data about real-world driver behavior, which is often private and sensitive.
Can we adjust MARL approaches to converge faster?
Although improvements in algorithm design or parameter tuning may improve convergence speed, our analysis (drawn from a broad spectrum of training configurations) suggests that we shall not expect to improve our results by a significant margin. Our trivial problem with a small action space and fleet size already required the equivalent of years of commute time to converge. Training algorithms for real-world tasks will converge orders of magnitude longer. Nevertheless, the following covers a selection of approaches proposed before our study which may improve the convergence rate and stabilize the learning process.
Ilhan et al. (2019) shows that off-policy learning using action advisement between agents can improve training efficiency. However, this approach has only been evaluated with 3 agents, which is insufficient for our study. Hong et al. (2017) shows that learning about other agents’ policies through observations can address non-stationarity. Yet, their method relies on the amount of information embedded in the agent’s raw observations, which is often limited. Alternatively, Wang et al. (2022) shows that if an agent learns how the strategies of other agents evolve with its behavior, defining a stability reward can stabilize learning. As promising as this notion is, it considers only pairwise stability, rendering it inapplicable for bigger systems.
What if reinforcement learning is not suitable for vehicle routing?
RL, as we discussed in Section 2.1, is already regarded as a viable solution for routing optimization, offering the ability to learn complex coordination and solve dynamic problems. Existing literature often frames this problem as a Traffic Assignment Problem (TAP), which seeks to satisfy a travel demand (number of travelers) navigating between origin-destination pairs within limited route capacities while minimizing costs, such as reducing system travel times. An approach to solve this problem includes methods based on the Frank-Wolfe (FW) algorithm (Frank et al., 1956), which is a general-purpose non-linear optimization method that is considered state-of-the-practice (Bar-Gera, 2002) but suffers from a slow convergence rate, as discussed in detail in Patriksson (2015).
Can AVs lead to more efficient urban traffic systems?
This study examines a future where AVs operate within today’s traffic network alongside human drivers. One could argue, in contrast to our position, that future cities will benefit from AV deployment. Specifically, Richter et al. (2022) show that AV deployment can offer advantages (including reduced congestion and shorter commutes), but in varying scales depending on the type of city. Makahleh et al. (2024) state that AV introduction can contribute to sustainability by improving air quality and reducing $NO_{2}$ concentrations. Lastly, Vahidi & Sciarretta (2018) mention the energy-efficient driving opportunities that CAVs can offer. However, Fagnant & Kockelman (2015) highlights the need for strict regulations in the deployment of AVs.
@@
Our position aims to shape the research focus on creating and developing algorithms that can optimize the collective routing of AVs. It demonstrates the potentially negative consequences for our future urban systems and their sustainability that are likely to occur if actions are not taken. We list gaps in the state-of-the-art that could render the large-scale deployment of AVs using MARL for urban routing decisions harmful to the overall system performance. To address these challenges, we propose research directions to leverage the capabilities of CAVs.
Appendix A Examples of real-world urban topologies corresponding to our two-route network
(a) Split, Croatia.
(b) Bochnia, Poland.
(c) Lisbon, Portugal.
(d) Tarnow, Poland.
Figure 7: Occurrence of two-route networks similar to the one presented in this study across cities worldwide (source: OpenStreetMap contributors (2017)).
Appendix B Experiments
The training for the experiments shown in Figure 4, Figure 5, Figure 6, and Figure 11 involves 200 policy updates and each update includes 40 frames collected per agent (episodes or days). The epochs are 10, the minibatch size is 2, and the activation function is the Tanh. In Figure 4 each algorithm experiment was conducted for ten independent runs, five runs in Figure 5 and Figure 6 and three runs in Figure 11. The error bars indicate the standard deviation across these experiments. For clarity, data points are recorded every 10 episodes and smoothed using the moving average method with a window size of value 10. Hyperparameter optimization was performed for each algorithm separately.

### v2

@@ 4 Problem statement
Figure 2: We model the routing "game" between humans and AVs on urban traffic networks using the RouteRL RouteRL2024 framework, which includes a custom PettingZoo terry2021pettingzoo environment, communicating with SUMO as it trains optimal routing policies with standard MARL algorithms via torchRL bou2023torchrl .
We illustrate our position on a simple Two-Route (Yield) network (TRY) (Figure 1a, b), which, while abstract, is carefully designed to capture arising issues of AV introduction in traffic networks. To assess the generalizability of our findings we also include a real-world case. In the TRY network, each agent selects from two possible routes (0 or 1) to reach the destination. Then, the reward (travel time) is collected from the environment to update choices for the next episode. Each episode is interpreted as a day on which the 22 drivers commute through the network. For clarity, the setting is unrealistically static. Each human driver follows the same route every day, departing at the same time, and the travel times do not vary day-to-day (none of the above holds in real networks as discussed in 2.3, see Appendix C). By design, the system has two Nash equilibria achievable by humans, one optimal (System Optimal - User Equilibrium, Figure 1a) and one suboptimal (Suboptimal - User equilibrium, Figure 1b). Humans mutate into AVs, which will use any MARL algorithms to find optimal policies. To simulate this scenario, we use the RouteRL RouteRL2024 framework (Figure 2), which models mixed-traffic scenarios, where AVs are simulated as RL agents and humans behave according to a given human-behavior model.
SUMO. An open-source, state-of-the-art, microscopic, agent-based traffic simulator used as the traffic environment in which each vehicle navigates the road network according to the IDM SUMO2018 .
Action. The action space of the agents corresponds to the set of available routes connecting their given origin with their destination and is discrete with value two on the TRY network (See 1a, b).
Reward. The reward $r_{i}^{t}(a_{i}^{t}|s_{i}^{t})$ is the negative travel time of each agent $i$ to reach from its origin to its destination, as calculated by SUMO.
Observation. We assume, plausibly for the future systems, that AVs observation $o_{i}^{t}$ is composed of their departure time and the number of agents that departed before them and chose routes 0 and 1.
Human agents. We follow the classical representation of human route-choice behavior from transportation research. Human drivers are rational decision-makers aiming to maximize their perceived utility Cascetta by selecting actions that minimize expected travel times. Their expectations are updated based on experienced travel times (from SUMO). In scenarios with adaptation, humans shift to an alternative route with $10\%$ probability.
@@
We consider two plausible equilibria resulting from human collective actions: first, when all humans select the shorter route ([$[\{0\}^{22}]$, 1a), and second when they all opt for the longer route ([$[\{1\}^{22}]$, 1b). Both meet Nash criteria for User Equilibrium Wardrop1952ROADPS (common paraphrase of Nash equilibrium for the route choice context), with the former being also System Optimal (minimizing total travel time in the system merchant1978optimality ). None of the drivers is inclined to change their route, as it would reduce individual rewards (travel longer and arrive later). For stability, we fix the route for the first three agents to regulate the early loading of the system. After 200 days of simulation, the system is stable (the next day the rational drivers will replicate today’s choices), fair (travel times are equal among drivers), and either globally optimal (Figure 1a where total travel costs are minimized) or suboptimal (Figure 1b). The individual and system costs (travel times) are reported in Table 1.
5 Experimental support for the position
@@
In the equilibrated human system, we first replace a random human vehicle with an RL-controlled AV. Its traffic properties (reaction, acceptance gap, acceleration profile, and other IDM parameters Treiber_2000 ) remain intact. AVs are indistinguishable from humans by all but routing decisions: they may use any RL (MARL) algorithm to converge to the optimal policy.
Figure 3: Single AV replicates human choices, and each tested RL algorithm finds the optimal solution to the binary choice problem (means and error bars computed across replications; see Section B).
We demonstrate that RL finds the optimal solution, and AV agents follow the crowd (replicate the decision of the human that drove that vehicle before). Unsurprisingly, since the problem is a trivial binary decision in a static environment, this is true for all suitable RL algorithms, including Deep Q-Learning (DQN), PPO, and SAC, as we show in Figure 3. In any equilibrated system, any RL algorithm training AVs with the same reward formulation, and action space as their human predecessors will replicate the optimal strategy, which can be derived directly from conditions of Nash equilibrium. Moreover, in a dense traffic environment, the impact of a single vehicle is unlikely to be noticed by other humans. The marginal cost of actions taken by a single AV to other humans will fall within what is known as the indifference band di2017indifference , making them indistinguishable from the traffic stochastic noise neun2023traffic4cast . Our position, that AVs will disequilibrate the traffic networks, does not hold for a single AV, that will converge to the solution of its human predecessor and will not impact other human drivers.
@@
Already in our simple scenario, serious issues arise when multiple vehicles learn optimal policies simultaneously. Most likely, more than one AV will be introduced into our cities (whenever they become ready to operate reliably), each solving the same problem of identifying optimal routing policies to reach its destination. Such a multi-agent setting allows a significant alteration to the initial problem: the AVs may communicate (becoming the so-called CAVs) and share information.
First, we demonstrate the natural first stage of AV introduction: with no communication and reward formulation identical to the one of the selfish human drivers. Now, the environment has become non-stationary jiang2024blackbox , as the state transitions and rewards of an agent are influenced by the evolving policies of other agents. With as few as 10 AVs (where non-convergence appears after the introduction of just 4 AV agents discussed in Section 5.4 and Appendix G, L) our setting is sufficient to argue for our position.
@@
Figure 4: (a, b): 8000 episodes (that correspond to over 20 years) of training for selected MARL algorithms. The trivial solution is found only by IPPO and MAPPO. Other algorithms fail to converge to the optimal policy. Convergence in the suboptimal equilibrium is achieved after 1000 days, but System optimal requires much more training, far exceeding the patience threshold for the remaining 12 human drivers in the system. The first 200 episodes represent the human learning phase, followed by the training and testing phases of the AVs. (c): The MAPPO solution (red), however, is not robust and falls to suboptimal as soon as non-determinism is introduced to the environment, demand, or human behavior (other lines). The line represents the mean with 0.15 and 0.85 percentiles. See Appendix A for the experiments’ setup.
In Figure 4(a, b), we present two arguments supporting our position. First is the class of algorithms that failed to converge after sufficiently long training (QMIX, VDN, IQL, MASAC, ISAC). Specifically, the choices are far from optimal, as reflected in the noisy profiles that fluctuate around 0.5. QMIX and VDN probably do not find optimal solutions due to the cooperative nature of their implementation. The difficulty of achieving stability and convergence in ISAC and MASAC can be attributed to their off-policy nature combined with the nonlinear function approximation used in neural networks DBLP:journals/corr/abs-1812-05905 ; NIPS2009_3a15c7d0 . The consequences of introducing MARL routing algorithms to AVs are negative to all parties: not only the travel times are longer for all agents (humans and AVs) but they also become variable, as we report in Table 1. IPPO and MAPPO, however, converged after a lengthy training. Theoretical explanations of the convergence difficulties are provided in Appendix E.
@@
Some MARL algorithms converged and successfully managed to identify optimal policies (the small joint action space of 1024 can be easily enumerated to explicitly identify optimal solutions). Eventually, the 10 AVs managed to stabilize their actions and reliably make optimal choices (Figure 4(a,b)). After that, the negative impact on the system and the humans diminishes (see Table 1). Nonetheless, the learning process was lengthy, with thousands of episodes (days).
The training can be interpreted twofold: within a virtual environment or in a real system. The former is neutral to the real system and its users since AVs can train their policies virtually and deploy them only after training is complete and the policies have converged to optimal solutions. Humans will not be affected and iterations remain only virtual. This, however, requires virtual environment suitable to train a policy, which, is not available to date, as we argued in Section 2.3. If so, the learning period needs to be treated physically and algorithmic iterations are not abstract episodes anymore, but physical days of real disturbances. Eventually, disturbances diminish (as for IPPO and MAPPO), yet the negative impact on the system and its users accumulates (to values presented in Section F). Alternatively, RL can be inaccurately trained on imperfect models and collected data and later fine-tuned in reality, as demonstrated in nair2021awacacceleratingonlinereinforcement . However, sim2real transfer presents significant challenges, as discussed in zhao2020sim . Real-world complexity exceeds any simulation’s capabilities.
@@
Figure 5: Introducing AVs on the Ingolstadt network (with 1000 agents and 1000 nodes) not only increased human travel times but introduced variability of travel times during training, which did not converge in 3000 days (MAPPO episodes). (see Appendix D).
As demonstrated, the problem behind our position is in multiple simultaneously learning agents. To find how many simultaneously learning agents can negatively affect the system, we simulated scenarios with a gradually increasing number of agents and reported when convergence issues arose. In each simulation, the specific AVs were chosen at random, with the condition that no two AVs are consecutive (there is always a human agent between AVs), and the first two vehicles never mutate. The algorithms start failing to converge even with 3-5 agents, as shown in Figure 8 and Appendix L. This reveals a systemic issue that becomes increasingly prevalent as the number of agents grows, which will pose a serious threat to traffic performance when more AVs participate in our daily commute.
@@
Another aspect of our position lies in the communication, collaboration, and/or centralization schwarting2019social - which makes AVs the CAVs. This allows us to better exploit the potential of autonomous driving, presumably at the cost of sharing private information with others or with central agents nayak2021autonomous . Can we trade our private destination and origin to resolve the non-stationarity issues? To some extent, centralization may speed up the convergence, yet nowhere close to solving the problem and leading to issues with combinatorically growing search space (see Figure 9 bottom in Appendix H).
5.6 Scaling from the abstract case to real-world scenarios - Ingolstadt results.
To reinforce our results, we reproduce our experiment and investigate what happens when complexity is added to the abstract case. We gradually introduce real-world phenomena (see Figure 4c and Table 2) from Section 2.3, successively including non-determinism in the traffic flow model (green), demand patterns (implemented as random departure times, yellow) and human adaptation (blue).
As these complexities are introduced, the system becomes increasingly disequilibrated. In response, humans will naturally seek ways to improve their payoffs (arrive faster) watling2013modelling . This behavior is similar to the process of finding the equilibrium bie2010stability , but even less predictable he2012modeling . This adds another source of non-stationarity to the system. We include a probabilistic adaptation formula (to the human decisions), and the previously optimal system now shifts to the suboptimal state (Figure 4c).
Table 2: MARL convergence under varying complexity levels (✓ = deterministic, ✗ = non-deterministic).
Network Traffic Flow Demand Adaptation Convergence
TRY yes
TRY no
TRY no
TRY no
Ingolstadt no
Finally, we report the results of the Ingolstadt network, using realistic demand, from the RESCO benchmark ault2021reinforcement (see Section D for more details). This demand includes 1000 agents with varying origins and destinations, each selecting between four paths, rendering the joint action space to $4^{1000}$. As shown in Figure 5, the average travel time of all the human agents in the system (pre-AV) is stable. However, after the introduction of AVs (episode 500), the previously stable system becomes unstable, exhibiting oscillatory travel times with values higher than before.
In all the above-mentioned cases, the central issue of our position remains present: agents did not converge to the optimal solution even after many episodes (4c), and during MARL training the system was destabilized (demonstrated as the variability of travel times on Figure 5).
@@ 6 Conclusion
AVs have begun appearing in traffic networks worldwide, and their presence is expected to increase in the coming years. As AVs become more prevalent, algorithms can be applied to optimize their route choices. In practice, the state-of-the-art MARL algorithms employed in this paper need hundreds of episodes to converge to optimal policies, even in trivial cases. The problem is amplified when more realistic traffic dynamics are introduced in the simulations. To realistically simulate human route-choice behavior, we miss theories, data to verify them, and simulation tools. Realistic urban mobility simulators are lacking, and the training episodes would probably need to be deployed directly on real traffic systems, disrupting the traffic networks. Needless to say, such a disruption should be avoided at all costs. We hope our contribution will spark discussions within the MARL community, encouraging the collaboration of authorities with the ML community to regulate autonomous collective routing. Such an experimental research program is needed to ensure we fully exploit the opportunities of the new technology (AVs) and algorithms to help us improve the traffic in future cities.
@@ 7 Alternative Views
Why regulate AV routing since it is not yet a problem? Safety issues are indeed the foremost concern in AV deployment, and it may well be more than a decade before the problems indicated become urgent. Nevertheless, if this happens when large fleets are allowed to group routes, it may completely disrupt the traffic that should be preempted and now is the right time to act.
Authorities should not monitor AI/ML-based routing algorithms. One could argue that authorities do not need to monitor AI/ML-based routing algorithms, since cities already manage infrastructure through traffic signals, congestion pricing, and physical designs. Furthermore, the competitive market incentivizes companies to develop efficient routing tools to offer superior services to their users. While these points are valid, history shows that markets may degenerate into vicious degrading competition if left alone, especially when novel strategies, whose efficiency may involve dumping, introducing chaos, etc. become available to group players.
What if car companies would not apply AI/ML for the routing policies of AVs? Artificial intelligence has already been integrated into many aspects of our daily lives. As discussed in Section 3, reinforcement learning has been applied to optimize vehicle routing. It is therefore natural to expect that such technologies will also be adopted by car companies.
What if reinforcement learning is not suitable for vehicle routing? RL, as discussed in Section 3, is already regarded as a viable solution for routing optimization, offering the ability to learn complex coordination and solve dynamic problems. However, this paper does not claim that MARL is the only method car companies might adopt. On the contrary, we argue that companies may deploy a range of algorithmic approaches—some of which could have even more detrimental effects on traffic flow. This is why such routing algorithms must be subject to monitoring.
What if the limitations of the virtual environment are specific to this study and could be resolved with better design or tools? As discussed in Section 2.3, accurately representing dynamic, multi-agent systems like traffic networks requires highly complex simulation environments. Modeling realistic traffic demand patterns involves acquiring and handling data about real-world driver behavior, which is often private and sensitive.
@@
Our position aims to direct the research focus towards assessing the algorithms that can optimize the collective routing of AVs. It demonstrates the potentially negative consequences for our future urban systems and their sustainability that are likely to occur if actions are not taken. We list gaps in the state-of-the-art that could render the large-scale deployment of AVs using MARL for urban routing decisions harmful to the overall system performance. To address these challenges, we propose research directions to leverage the capabilities of CAVs. We believe it is time to act now before the damage has been done.
Appendix A Experiments
The experiments were conducted using version 0.0.1 of the RouteRL framework [3] for the TRY network and version 1.0.0 for the Ingolstadt network. RouteRL is released under the MIT License. All experiments on the TRY network involve 200 policy updates, with each update consisting of 40 frames collected per agent (interpreted as episodes or days). Training is performed for 10 epochs, using a minibatch size of 2, and the Tanh activation function. The algorithm scripts are derived from the state-of-the-art (SOTA) implementations provided by the TorchRL library [10]. All experiments were conducted using version 0.3.0 of the open-source TorchRL library to ensure reproducibility.
In the experiment conducted on the Ingolstadt network, 4,000 frames were used with 800 policy updates. The AVs were introduced into the traffic network from episode 500. The training was performed using 1 epoch and a minibatch size of 32. All hyperparameters were consistent with those listed in Table 3.
Each experiment was replicated a different number of times, depending on the scenario. In the case of multiple simultaneously routing AVs (described in Section 5.2 and shown in Figure 4(a, b)), each algorithm experiment had 10 replications for each equilibrium. For the experiments investigating the minimum number of simultaneously routing AVs that can destabilize the traffic network (described in Section 5.4 and shown in Figure 8), each algorithm was replicated for five independent runs for each equilibrium. The experiments exploring the effects of non-determinism and centralization (described in Sections 5.6, 5.5 and shown in Figure 4c and 9) were conducted for 3 independent replications per setting. Hyperparameter optimization was performed for each algorithm and setting (e.g., centralization) separately.